A proton (mass m, charge +e) and an alpha particle (mass 4m, charge +2e) approach one another with the same initial speed v from an initially large distance. How close will these two particles get to one another before turning around?
1
Expert's answer
2020-05-18T10:17:23-0400

The kinetic energy of the moving particles is completely transformed into electric potential energy at the point of closest approach.

"U_e \t = \tK\\\\ \t \n \n\\frac{k(e)(2e)}{r^2} \t\n = \t\\frac12 \t (m)v^2 + \t\\frac12 \t (4m)v^2"


"r=\frac{4ke^2}{5mv^2}"